solution a=50%, solution b =30% what% of each do i need to get 45%
Answer by rwm(914)

.5a+.3b=.45(a+b)
0.50a+0.30b = 0.45a+0.45b
0.05a-0.15b = 0
a=3b
b=a/3
b= 1/3*a